Calculate 1 percent of half of 5 billion divided by 5
Found 2 solutions by erica65404, stanbon:
Answer by erica65404(394)

Half of 5 billion
2.5 billion
1% of 2.5 billion
2.5*.01=25,000
25,000/5=5000
The answer is 5000

Answer by stanbon(75887)

Calculate 1 percent of half of 5 billion divided by 5
------
(1/100)*(1/2)(5 billion)/5
= 1 billion/200
= 1,000,000,000/200
= 10,000,000/2
= 5,000,000
-----------------
Cheers,
Stan H.